four balls are there

i pick up two balls out of them

two of them are white

what is the probability that all of them are white?

Grade:12

5 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

Required probabilty=2/4=1/2

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

vikas askiitian expert
13 years ago

total balls = 4

white balls = 2

probablity that  white ball is taken out = 2/4= 1/2

now only 1 white ball is remaining

total balls = 3

probablity that white ball is taken out = 1/3

total probablity = multiplication of both = 1/6

1/6 is the required ans

Mir Noaman Ali IIT-Roorkee
13 years ago

probability = favourable /total cases   in this case favourable case is choosing 2 white out of 4 white nd total is choosing 2 white out of 2white nd 2 of sum oder colour + choosing 2 white out of 3white and 1 of sum oder colour nd choosing 2 white out of 4 white hence 4c2 / (4c2 + 3c2 + 2c2) calculating it we get 6/11  approve the answer if u have understood the solution 

Mir Noaman Ali IIT-Roorkee
13 years ago

probability = favourable /total cases   in this case favourable case is choosing 2 white out of 4 white nd total is choosing 2 white out of 2white nd 2 of sum oder colour + choosing 2 white out of 3white and 1 of sum oder colour nd choosing 2 white out of 4 white hence 4c2 / (4c2 + 3c2 + 2c2) calculating it we get 6/10 = 3/5   approve the answer if u have understood the solution ..............

Sudheesh Singanamalla
13 years ago

Number of balls = 4

number of balls picked = 2

colour of 2 balls picked = white

the probability that all of them are white is 1

because

1. all 4 could be white

2. 3 white , 1 of another colour

3. 2 white and 2 of one colour each

the number of white balls could be either 2 or 3 or all 4 of them !

so probabilty that all of them are white is 1/2 or 3/4 or 1

since we get 2 white balls out of 2 balls we pick , the probabilty is 2/2 = 1

 

Please approve